# On some indices of foliations and applications

### v1

@@ 7.4. Alcántara–Mozo-Fernández’s conjecture
To continue, we present an example that answer the conjecture posed by Alcántara–Mozo-Fernández.
Example 7.7.
Let $\zeta,\lambda_{1},\lambda_{2}\in\mathbb{C}^{*}$ and $n\in\mathbb{N}$ with $n\geq 2$ such that $\lambda_{1}+\lambda_{2}(n+1)=0$. Consider the foliation $\mathcal{F}$ in homogeneous coordinates $[x:y:z]\in\mathbb{P}_{\mathbb{C}}^{2}$ given by
$\Omega_{\zeta,\lambda_{1},\lambda_{2}}=-\zeta(n+1)\lambda_{2}yx^{n}dx+((\lambda_{1}+\lambda_{2}n)zy^{n}-\zeta\lambda_{1}x^{n+1})dy+\lambda_{2}y^{n+1}dz.$
Then we have the following properties:
  1. (1)
    $\mathcal{F}$ is a non-dicritical logarithmic foliation on $\mathbb{P}_{\mathbb{C}}^{2}$ has degree $n$ with a unique singularity $[0:0:1]$.
  2. (2)
    The curve $C=\{y(y^{n}z-\zeta\cdot x^{n+1})=0\}\subset\mathbb{P}_{\mathbb{C}}^{2}$ is quasi-homogeneous, $\mathcal{F}$-invariant and $[0:0:1]\in C$.
In fact, let $A(x,y,z)=-\zeta(n+1)\lambda_{2}yx^{n}$, $B(x,y,z)=(\lambda_{1}+\lambda_{2}n)zy^{n}-\zeta\lambda_{1}x^{n+1}$, and $C(x,y,z)=\lambda_{2}y^{n+1}$. Therefore,
$\Omega_{\zeta,\lambda_{1},\lambda_{2}}=Adx+Bdy+Cdz\quad\text{and}\quad Ax+By+Cz=0.$
Since the degree of $A$, $B$ and $C$ is $n+1$, we get $\Omega_{\zeta,\lambda_{1},\lambda_{2}}$ defines a foliation $\mathcal{F}$ of $\mathbb{P}^{2}_{\mathbb{C}}$ of degree $n$. On the other hand, a direct calculation shows
(30) $\Omega_{\zeta,\lambda_{1},\lambda_{2}}=\lambda_{1}\frac{dy}{y}+\lambda_{2}\frac{d(y^{n}z-\zeta\cdot x^{n+1})}{y^{n}z-\zeta\cdot x^{n+1}}.$
Since $\lambda_{1}+\lambda_{2}(n+1)=0$, we deduce that $\mathcal{F}$ is a logarithmic foliation. Analyzing the chart $z=1$ of $\mathbb{P}^{2}_{\mathbb{C}}$, we have $\mathcal{F}$ is given by
$\omega=-\zeta(n+1)\lambda_{2}yx^{n}dx+((\lambda_{1}+\lambda_{2}n)y^{n}-\zeta\lambda_{1}x^{n+1})dy.$
We get $p=[0:0:1]$ is a singularity of $\mathcal{F}$ with $\mu_{p}(\mathcal{F})=n^{2}+n+1$, implying that $p$ is the unique singularity of $\mathcal{F}$. By (30), we obtain that $C$ is $\mathcal{F}$-invariant, moreover the germ of $C$ at $p$ is a quasi-homogeneous curve with $\mu_{p}(C)=n^{2}+n+1=\mu_{p}(\mathcal{F})$. Finally, to prove that $p$ is a non-dicritical singularity of $\mathcal{F}$ is suffices to find the a local expression to the strict transform of $\mathcal{F}$ by a blow-up at $p$. Of course, taking local coordinates $\pi(x,t)=(x,tx)=(x,y)$ of the blow-up of $\mathbb{C}^{2}$, we get
$\tilde{\omega}=\frac{\pi^{*}(\omega)}{x^{n}}=(-\zeta(n+1)\lambda_{2}xt+t((\lambda_{1}+\lambda_{2}n)t^{n}-\zeta\lambda_{1}x))dx+x((\lambda_{1}+\lambda_{2}n)t^{n}-\zeta\lambda_{1}x)dt$
The strict transform $\tilde{\mathcal{F}}$ of $\mathcal{F}$ by $\pi$ is defined by $\tilde{\omega}$ and since the exceptional divisor $\{x=0\}$ is $\tilde{\mathcal{F}}$-invariant, we conclude that $p$ is a non-dicritical singularity of $\mathcal{F}$. These verify the above items.

### v2

@@ 7.4. Alcántara–Mozo-Fernández’s conjecture
Now, we present an answer to the conjecture posed by Alcántara and Mozo-Fernández in the case where $C$ is reduced.
Theorem 7.7.
There does not exist a logarithmic, non-dicritical, quasi-homogeneous foliation $\mathcal{F}$ on $\mathbb{P}^{2}_{\mathbb{C}}$ of degree $d$, with a unique singular point $p$, that admits an invariant reduced algebraic curve $C$ of degree $d_{0}\geq 2$, containing all separatrices passing through $p$.
Proof.
Suppose, by contradiction, that there exists a logarithmic, non-dicritical, quasi-homogeneous foliation $\mathcal{F}$ on $\mathbb{P}^{2}_{\mathbb{C}}$ of degree $d$, with a unique singular point $p$, and an invariant reduced algebraic curve $C$ of degree $d_{0}\geq 2$ which contains all the separatrices passing through $p$.
We fix the singular point at $p=[0:0:1]$, and consider local coordinates $(x,y)\in\mathbb{C}^{2}$ centered at $p$. Since $C$ contains all the separatrices passing through $p$, and $\mathcal{F}$ is non-dicritical at $p$, it follows that $C$ is the total union of the separatrices of $\mathcal{F}$ at $p$. Moreover, if the germ of $C$ at $p$ is given by $C_{p}=\{f(x,y)=0\}$, then, by hypothesis, $f$ is quasi-homogeneous. Hence, after a suitable change of variables, we can write
$f(x,y)=u\cdot x^{m}\cdot y^{n}\Pi_{j=1}^{s}(y^{k}-\zeta_{j}x^{q}),$
where $u$ is a unit, $m,n\in\{0,1\}$, $\gcd(k,q)=1$, and $\zeta_{j}\in\mathbb{C}^{*}$ for all $j=1,\ldots,s$. Without loss of generality, we may assume that $k<q$. Now, we analyze the possible branches of $f$. First, note that $m=0$, because the intersection $\{x=0\}\cap\{y^{k}z^{q-k}-\zeta_{j}x^{q}=0\}$ consists of two points, namely $[0:0:1]$ and $[0:1:0]$, which would yield two singularities of $\mathcal{F}$, contradicting our hypothesis. Additionally, $C$ cannot have two different branches of the form $C_{j}=y^{k}-\zeta_{j}x^{q}$; otherwise, taking two such branches, say $C_{i}=y^{k}-\zeta_{i}x^{q}$ and $C_{j}=y^{k}-\zeta_{j}x^{q}$, we obtain, by Bézout’s theorem
$I_{p}(C_{i},C_{j})=kq=q^{2},$
which implies that $k=q$, yielding a contradiction. Consequently, we conclude that $f(x,y)=uy(y^{k}-\zeta x^{q})$, where $\zeta\in\mathbb{C}^{*}$. In particular, $d_{0}=q+1$, and since $\mathcal{F}$ is logarithmic, $d=d_{0}-2=q-1$ by (27). Moreover, since $\mathcal{F}$ is a non-dicritical generalized curve foliation, we obtain by [7, p. 19] and [9, Theorem 4]
$d^{2}+d+1=\mu_{p}(\mathcal{F})=\mu_{p}(C)=\mu_{p}(y)+\mu_{p}(y^{k}-\zeta x^{q})+2q-1,$
which implies that $q=k+1$. Thus, we conclude that $f(x,y)=uy(y^{k}-\zeta x^{k+1})$. In particular, $f$ has two branches and contains all separatrices of $\mathcal{F}$ passing through $p$. Now, observing that in homogeneous coordinates $[x:y:z]$ of $\mathbb{P}^{2}_{\mathbb{C}}$, the curve $C=\{y(y^{k}z-\zeta x^{k+1})=0\}$ is an algebraic curve invariant by $\mathcal{F}$, and since $d_{0}=q+1=k+2$ and d=q1=kd=q-1=k, we conclude that $\mathcal{F}$ can be defined by a closed logarithmic 1-form with poles along $C$, by Cerveau’s Theorem [11, Theorem 1.4]. That is, there exist $\lambda_{1},\lambda_{2}\in\mathbb{C}^{*}$ satisfying $\lambda_{1}+(k+1)\lambda_{2}=0$, and
$\mathcal{F}:\quad\Omega=\lambda_{1}\frac{dy}{y}+\lambda_{2}\frac{d(y^{k}z-\zeta x^{k+1})}{y^{k}z-\zeta x^{k+1}}.$
However, since $\lambda_{1}+(k+1)\lambda_{2}=0$, we can write $\Omega$ as:
$\Omega=\lambda_{2}\frac{y^{k+1}}{(y^{k}z-\zeta x^{k+1})}d\left(\frac{y^{k}z-\zeta x^{k+1}}{y^{k+1}}\right),$
which implies that $\mathcal{F}$ admits a rational first integral, and consequently, has a unique dicritical singularity at $p=[0:0:1]$. This contradicts the assumption that $\mathcal{F}$ is a non-dicritical foliation at $p$. ∎
Acknowledgments
The authors are grateful to Marcelo E. Hernandes and Hernán Neciosup Puican for their remarks.
They would also like to thank the anonymous referees whose comments on a previous version of this article helped to improve both the content and the presentation.
Funding information
The authors gratefully acknowledge the support of Universidad de La Laguna (Tenerife, Spain), where part of this work was done: Spanish grants PID2019-105896GB-I00 funded by MCIN/AEI/10.13039/501100011033 and PID2023-149508NB-I00 funded by MICIU/AEI /10.13039/501100011033 and by FEDER, UE. This work was also funded by the Dirección de Fomento de la Investigación at the PUCP through grant DFI-2024-PI1100. The first author acknowledges support from CNPq Projeto Universal 408687/2023-1 ”Geometria das Equações Diferenciais Algébricas” and CNPq-Brazil PQ-306011/2023-9.
Data availability
We do not analyze or generate any datasets, because our work proceeds within a theoretical approach.
Conflict of interest
The authors declare that they have no conflict of interest.